### v1

# An Industry Interview Study of Software Signing for Supply Chain Security

Abstract
Many software products are composed by the recursive integration of components from other teams or external parties. Each additional link in a software product’s supply chain increases the risk of the injection of malicious behavior. To improve supply chain provenance, many cybersecurity frameworks, standards, and regulations recommend the use of software signing. However, recent surveys and measurement studies have found that the adoption rate and quality of software signatures are low. These findings raise questions about the practical application of software signing, the human factors influencing its adoption, and the challenges faced during its implementation. We lack in-depth industry perspectives on the challenges and practices of software signing.
To understand software signing in practice, we interviewed 18 high-ranking industry practitioners across 13 organizations. We provide possible impacts of experienced software supply chain failures, security standards, and regulations on software signing adoption. We also study the challenges that affect an effective software signing implementation. To summarize our findings: (1) We present a refined model of the software supply chain factory model highlighting practitioner’s signing practices; (2) We highlight the different challenges – Technical, Organizational, and Human — that hamper software signing implementation; (3) We report that expert subjects disagree on the importance of signing; (4) We describe how failure incidents and industry standards affect the adoption of software signing and other security techniques. Our findings contribute to the understanding of software supply chain security by highlighting the impact of human and organizational factors on Software Supply Chain risks and providing nuanced insights for effectively implementing Software Supply Chain security controls — towards Software signing in practice.
1 Introduction
Like all engineered products, software can be built more quickly when components are reused from other sources [1, 2]. This integration accelerates production but requires trusting external parties [3]. This trust increases when the provenance of components is known [4]. The strongest guarantee of provenance is a cryptographic signature created using a secret known only to the vendor. This technique, known as software signing, is underutilized by many software producers [5, 6]. However, as a result of software supply chain attacks such as SolarWinds, signing has become integral to many government, academic, and industry proposals to secure software supply chains [7, 8, 9].
Although software signing has desirable properties, it must be adopted carefully to bring benefits. At present, measurement studies have shown that missing or erroneous signatures are common across the software supply chain [5, 10, 11, 6, 12]. When signing is omitted or done unsuccessfully, researchers have demonstrated exploits [8, 13, 14, 15]. Existing security frameworks recommend software signing, but the ambiguous guidelines across these frameworks lead to a lack of a detailed implementation model [16, 17]. In order to adopt software signing, organizations would benefit from a detailed understanding of industry practices and challenges. The academic and grey literature does not offer detailed industry perspectives on the experiences that engineers have when adopting software signing.
Our work represents the first qualitative study of software signing in industry practice. We interviewed 18 industry security experts and decision-makers across 13 organizations. Our interviews focused on four aspects: the software signing practices employed in industry; the importance accorded to software signing in practice; the influence of industry standards and failure events; and the organizational, technical, and other challenges that hamper the effective implementation of software signing in practice.
We present a comprehensive summary of our findings. We refined the software supply chain factory model to show the roles of signing. While artifacts are frequently signed, signature verification is notably low. Our study reveals that the most common challenges in practice are technical, organizational, and human factors. These include key management and infrastructure, compatibility, a lack of signature verification, expertise in setting up signing tools, and developer attitudes. Our research indicates that software signing is typically used in one of three ways to mitigate risks: establishing provenance, serving as an auxiliary security technique, or fulfilling requirements akin to ticking a checkbox. While software supply chain failures, regulations, and other industry standards do influence the adoption of software signing and other security techniques, our participants have identified several issues associated with these regulations and standards.
Our contributions are:
  1. 1.
    We refined the software supply chain factory model to reflect diverse industry practices and the forms that signing may take therein (Figures 4 and III).
  2. 2.
    We identified common technical, human, and organizational challenges in adopting signing (Table IV).
  3. 3.
    We report on the relative importance of signing in software supply chain security strategies (section 5.3).
  4. 4.
    We discuss the effect that failures, standards, and regulations have on software signing practices (section 5.4).
Significance: Software signing is a widely recommended technique, and has become mandatory for US federal software vendors. Many organizations are therefore seeking to integrate software signing into their engineering processes. This study is the first qualitative work on software signing in industry. We provide insights on software signing in industry: adoption practices, challenges, and human and organizational factors to consider. Such knowledge will guide organizations in their software signing journey, and contribute to the software engineering community’s strategies for enhancing software supply chain security.
2 Background & Related Works
Here, we cover software supply chains (section 2.1), cybersecurity considerations (section 2.2), and empirical studies (section 2.3).
2.1 Software Supply Chains
Software production is often depicted using variations on the early Software Factory Model [18, 19] or the more recent Software Supply Chain Factory Model (SSCFM) [20, 21], depicted in Figure 1. The SSCFM approximates software product from the perspective of one software engineering organization or team. The SSCFM model shows that a software product integrates first-party and third-party components, that these components are packaged for downstream use, and that the software product is at risk of malicious code injection from internal or external parties.
Figure 1: The Software Factory Model [20].
The study of software supply chains has been driven by engineering organizations’ greater reliance on external components. Many works have reported on this in various software contexts. For example, a 2024 Synopsys report stated that, in 1,067 projects across 17 industry sectors, $\sim$96integrated open-source components, and $\sim$77of total code is from open source [22]. Even in government- and safety-critical systems, the use of third-party open-source components has moved from anathema to normal [23, 24]. Analyses of open-source software package registries have shown large increases year-over-year in the number of available components, the number of actively-used components, and the number of transitive dependencies [25, 26, 27, 28].
When an external component contains a vulnerability, that vulnerability may be propagated to its dependents. The increased reliance on external components in software production has amplified the potential for exploitation of these vulnerabilities, and many attack scenarios have been proposed [29, 30, 31, 32, 33]. For example, Lauinger et al. [34] report that $\sim$37of the most popular Internet websites rely on at least one library with a publicly acknowledged vulnerability. Stock et al. showed that among 1,273 real-world client-side XSS vulnerabilities, 273 cases occurred exclusively within third-party code [35]. Upstream in the supply chain, Zimmerman et al.’s analysis of npm indicated that up to 40% of all packages depended on code with at least one publicly known vulnerability [26]. Given the prevalence of such inherited vulnerabilities, Zahan et al. [36] and similar studies have summarized attack vectors and potential weak links in software supply chains [37, 38, 39, 40, 41, 42, 43, 44].
2.2 Securing Software Supply Chains
2.2.1 Methods and Practices
Numerous methods and tools have emerged to improve software supply chain security. Okafor et al. [45] characterize them as promoting three distinct dimensions: transparency (knowing all components in the chain), separation (isolating components), and validity (ensuring integrity of components). For example, transparency is increased by adoption of Software Bills of Materials (SBOMs) [46, 47]. Separation is promoted by techniques such as sandboxing [48, 49, 50, 51, 52] and zero-trust architectures [53, 54]. Validity is improved by reproducible builds [55, 56], software signatures [5], and attestations about artifact metadata such as authorship [57].
Security standards, frameworks, and regulations have been established to complement and direct these security efforts. Notable examples include The Linux Foundation’s Supply Chain Levels for Software Artifacts (SLSA) [20], the Cloud Native Computing Foundation’s (CNCF) Software Supply Chain Best Practices [58], Microsoft’s Supply Chain Integrity Model (SCIM) [59], and the US National Institute of Standards and Technology’s (NIST) Secure Software Development Framework (NIST-SSDF) [60] and NIST 800-204D [61], which were developed in conjunction with US Executive Order 14028 (EO-14028) [62]. At present, national regulations are inadequate to force compliance with secure supply chain practices [63].
2.2.2 Software Signing in Detail
Software signing is a formally guaranteed method of establishing the authorship of a software component [64, 65, 66, 14]. Signing ensures the provenance of components through public-key cryptography (Figure 2), improving validity (provenance) in supply chains [45]. Many standards and government regulations recommend signing as a base Software Supply Chain security technique [20, 7, 67, 59, 9]. For example, in The Linux Foundation’s SLSA framework, software signing is needed for Build Security Level 3 [20].
Figure 2: Typical workflow for software signing and verifying signatures. The component author packages (A) and signs (B) their software. The signed package (C) and public key (D) are published. To use a package, a user downloads it (E) and its public key (F) and verifies the signature (G).
2.3 Empirical Work on SW Supply Chain Security
The theoretical properties of techniques to improve software supply chain security are well known, whether via transparency, separation, or validity. However, adopting them in practice remains a challenge. Empirical/in vivo studies on software supply chain security have examined adoption practices and challenges to identify gaps in theoretical guarantees or usability. Some studies examine the general activities of component management, e.g., studies on dependency selection [68] and maintenance [69], and examinations of security challenges and incident handling procedures [70]. Others have examined specific techniques for supply chain security. For example, with respect to the transparency property, researchers have studied the practices of SBOM adoption [71, 72], and reported gaps in tooling and interoperability as well as incomplete adoption and maintenance across the software development lifecycle.
With respect to software signing, researchers have examined gaps in usability and adoption. Older implementations of signing were complex to use [73, 74], inspiring newer approaches with simplified tooling and support for key management [75, 76, 77]. About 25% of publicly-reported attacks from 2010-2021 exploited the missing or erroneous use of signing [15]. Gokkaya et al. showed that the top signing-related attack vectors were self-signed code, broken signature systems, and stolen certificates [42]. Several reports show that signing adoption rates for open-source components are low. Zahan et al. found that $\leq$0.5of packages in the npm and PyPI registries had signed releases [6]. Schorlemmer et al. corroborated this proportion, further noting that signing adoption is correlated with registry policies and that the correctness or quality of signatures depends on usable tooling [5]. Ladisa et al.’s survey shed light on these measurements, finding that software signing is valued but is considered moderately costly to set up [44, 78].
3 Knowledge Gaps & Research Questions
In our analysis of the literature discussed in section 2.3, we observed that our empirical understanding of software signing challenges and adoption is primarily derived from open-source software components. The available data on industry practices are surveys [44, 78]. We lack in-depth qualitative descriptions of industry experiences, creating a gap in our understanding of software supply chain security. We do not know which signing strategies are effective, the challenges associated with implementing these methods, the roles played by suggested methods and tools, and the factors influencing their adoption. To fill this gap, we ask four research questions:
  • RQ1:
    How and why is software signing implemented by software teams?
  • RQ2:
    What are the challenges that affect the implementation and use of software signing?
  • RQ3:
    What is the perceived importance of Software Signing in mitigating risks?
  • RQ4:
    How does software supply chain failures and standards influence the adoption of software signing?
4 Methodology
We conducted semi-structured interviews with subjects who were cybersecurity practitioners. We depict our methodology in Figure 3. In this section we discuss our research design rationale (section 4.1), protocol development (section 4.2), participant recruitment (section 4.3, data analysis (section 4.4), and study limitations (section 4.5).
Figure 3: Study methodology. Several sources (left) informed our protocol topics (center). We applied thematic and framework analyses (right).
4.1 Approach — Research Design Rationale
Our questions require insight into practitioner perspectives. We considered two factors in selecting a methodology. First, since the prior work in this space is limited, there was relatively little basis on which to structure a closed-ended survey. Second, we expected recruitment to be a problem, since the target population has high security expertise and a broad organizational view, and thus are likely to be high-ranking (and busy) people. We therefore opted for a semi-structured interview [79], which would permit us to capture emergent results [80] (factor 1), and to leverage our professional relationships to improve response rates (factor 2).
The resulting data is expected to be exploratory, and rooted in practitioners’ perspectives and experiences [79]. The phenomenon of interest is software signing as a security method, which is a contextual decision based on an organization’s distinct policies, security expectations, perceived threats, and customer requirements [81]. We therefore adopt an interpretivist perspective for our research methodology design and analysis. An interpretivist approach treats each situation as unique [82], focusing on in-depth variables and factors related to each context [83].
4.2 Protocol Design and Development
To create the initial interview protocol, we studied three kinds of resources. First, to ensure a common language and a comprehensive set of topics, we studied software supply chain security frameworks such as SLSA, CNCF, SCIM, and NIST-SSDF (section 2.2). These frameworks use variants of the Software Supply Chain Factory Model (SSCFM, cf. Figure 1). We therefore used the SSCFM as our reference framework for structuring the C Sections of our interview protocol. This choice allowed us to investigate signing practices at various stages of the SDLC while using a model familiar to our subjects. Second, we examined grey literature sources to gain insights into some practical uses of software signing. Our primary reference was the official Sigstore blog as of Jan. 31, 2023. This resource has eight user reports about adopting software signing in various industry contexts, from which we derived questions about the contextual and organizational factors that subjects consider. Third, we reviewed adjacent literature [68, 69, 84] to understand effective questions related to software supply chain security, while avoiding duplication.
After protocol development, we refined the protocol through practice interviews (2) and pilot interviews (2). For practice, the lead author interviewed two of the secondary authors. These practice sessions led to modifications in sections B and C of our interview protocol (see Appendix Appendix C). After recruiting subjects, we treated the first two subjects as pilot interviews. During these pilots, we gathered additional insights into how subjects perceived software supply chain attacks and reorganized the protocol, particularly focusing on questions related to the perceived sources of greatest risk (section B). Since few questions were changed or added, we included the data from the pilot interviews in our study where feasible. The final change occurred after the seventh interview. The author team discussed the results so far and identified one emergent topic: how practitioners’ organizations integrate feedback into their SSC security processes. One question was added for use in the remaining interviews.
Our interview protocol is summarized in Table I, with a full version in Appendix Appendix B.
TABLE I: Summary of interview protocol. See Appendix Appendix B for full protocol. Topic D was not analyzed in this study.
Topic (# Questions) Sample Questions
A. Demographics (4) What is your role in your team?
B. Perceived software supply chain risks (4) Can you describe any specific software supply chain attacks (e.g., incidents with 3rd-party dependencies, code contributors, OSS) your team has encountered?
C. Mitigating risks with signing (8) How does the team use software signing to protect its source code (what parts of the process is signing required)?
D. Signing tool adoption (5) What factors did the team consider before adopting [TOOL/METHOD] over others?
4.3 Data Collection
Our Institutional Review Board (IRB) approved this study.
TABLE II: Subject Demographics. For anonymity, we used generic job roles, not specific titles. “Senior management” refers to senior managers, directors, and executives. “Technical leader” — senior, lead, partner, and principal engineers. “Engineer” and “Manager” — junior staff.
ID Role Experience Product Area
S1 Research leader 5 years Digital technology
S2 Senior mgmt. 15 years SSC security
S3 Senior mgmt. 13 years SSC security
S4 Technical leader 20 years Cloud, Dev tools
S5 Engineer 2 years Cloud security
S6 Technical leader 27 years Digital technology
S7 Manager 6 years Cloud security
S8 Technical leader 8 years Internet services
S9 Engineer 2.5 years SSC security
S10 Engineer 13 years Dev tools
S11 Technical leader 16 years Social technology
S12 Technical leader 4 years Cloud, OSS security
S13 Senior mgmt. 16 years Telecommunications
S14 Research leader 13 years Cloud security
S15 Senior mgmt. 15 years Aerospace security
S16 Technical leader 26 years Digital technology
S17 Senior mgmt. 15 years SSC security
S18 Manager 11 years Internet services
Target Population: To address our research questions effectively, the target population was industry practitioners with expertise in software signing. An ideal subject would have in-depth knowledge and experience related to software supply chain practices within their organizations. Such experts would have high credibility, insider perspectives, and nuanced interpretations.
Recruitment: We employed a non-probability-based Purposive and Snowball population sampling approach to recruit such subjects [82]. Our initial subject sample pool was members from the Cloud Native Computing Foundation’s Kubecon 2023 conference organizing committee, many of whom also contribute to software supply chain security projects hosted by The Linux Foundation. This pool contained many security experts, although we acknowledge that the connection to specific signing projects may bias their responses to this genre of signing approach. Subsequently, we expanded our subject pool through recommendations from this initial group of subjects and other industry contacts. We offered each participant a \$100 gift card as an incentive.
In total, we contacted 30 candidates, of whom 18 agreed to participate (60% response rate). These subjects came from 13 different organizations. Their demographic information is summarized in Table II. Subjects were affiliated with companies with security products or were part of their company’s security team. For subject anonymity, we do not distinguish those from CNCF or from snowballing.
Interviews: Interviews were conducted by the lead author via Zoom. The median interview duration was 50 minutes.
4.4 Data Analysis
Interview recordings were transcribed by www.rev.com (human transcription). We anonymized identifying information. Our analysis then proceeded in several stages.
(Stage 1) We began with data familiarization [85]. Three analysts (A1–A3) undertook an inductive reading, memoing, and coding process for all transcripts.
(Stage 2) Subsequently, following O’Connor & Joffe’s recommendation [86] to initially independently code 10-25% of the data, we randomly selected approximately 33% (6 transcripts) for multiple memoing. Two analysts (A1, A2) conducted this multiple memoing of data in three rounds, collaborating to discuss emerging code categorizations after each round. This process yielded 506 coded memos, which were reviewed by Analysts A1 and A3 over four one-hour meetings to create a codebook comprising 36 code categories, excluding demographics.
(Stage 3) To ensure the reliability of our coding, analysts A1 and A3 randomly selected and coded a transcript. We used the Percentage agreement metric as recommended by Feng [87] for easy coding tasks with nominal data. Since our codebook was collectively defined by both analysts A1 and A3, we assumed this coding task should not be too difficult — we achieved an 89% agreement rate, confirming this assumption. Any discrepancies in coding were resolved through discussion and code redefinition.
(Stage 4) Next, a second set of 6 transcripts was selected and coded by Analyst A1 using an initial codebook, adding 6 minor categories, each with less than 6 codes. This updated codebook was reviewed by Analyst A3.
(Stage 5) The same process was applied to the final 6 transcripts, yielding 4 more minor categories with fewer than 4 codes each. Out of the 10 new categories, only 2 provided new insights into our data, as determined through multiple review rounds by Analysts A1 and A3.
(Thematic and Framework Analyses) From this process we induced themes following Braune & Clark’s method [88]. These themes helped us answer RQs 2-4 directly. To answer RQ1, we also applied a framework analysis [89], mapping the themes to the software supply chain factory model as a reference framework to analyze participants’ responses and further categorize the themes.
Through our coding phase’s thematic analysis, we identified eight themes. Four of these themes were focused on the practical implementation of Signing, while the remaining four addressed various aspects of software signing: the challenges encountered, its perceived importance, the impact of security failures on its adoption, and the influence of security standards and regulations on its adoption.
To determine whether additional aspects of the phenomenon might be expected to emerge if we continued to sample, we measured saturation in the 18 interview transcripts. Following Guest et al.’s [90] recommendation, we tracked the cumulative appearance of new codes in each interview. Each interview had a median of 33 unique codes, also depicted in  Figure 5. Saturation was achieved after interview 11 (details in Appendix Appendix D).
4.5 Study Limitations
Positionality Statement: We acknowledge the role that our backgrounds may have played in this study [91]. The author team are cybersecurity researchers with expertise in software signing, so our expertise both enabled and influenced the protocol development and analysis. Beyond this, one of the authors is a contributor to a popular signing tool. To reduce bias in our study, that author had no involvement in the development of the research protocol or analysis. However, that author reviewed the results and framing of the study, to increase its validity for the software signing community.
Other Threats to Validity: This work is qualitative, and bears the limitations of any qualitative study. First, the cost of data collection reduces the sample size, threatening generalizability. Our subject recruitment may also have introduced biases, e.g., towards CNCF/The Linux Foundation-oriented frameworks and tools. As mitigations, we note that our codebook saturated at the 11th interview, with few/no new codes observed in subjects 12-18 (Appendix D); and our sample size is typical for similar studies [71, 72]. Second, qualitative data implies subjective analysis. To improve the reliability of our results, we followed an iterative process to develop the codebook and used multiple raters with a high measure of agreement.
Specific to the topic of software signing, the frameworks for software supply chain security and the available automation for implementing signing are evolving. Our results should be understood as a snapshot of current practices in industry. They will grow out of date as signing standards, regulations, and tooling continues to evolve.
5 Results
5.1 RQ1: Signing Implementations in Practice
Finding 1. We refine the software supply chain factory model from Figure 1, as shown in Figure 4 and Table III. Our refined model shows candidate points for the establishment of provenance. We analyze our subjects’ responses using this refined model, identifying points on this model where subjects’ teams or organizations employ software signing.
Figure 4: Our refined software supply chain factory model (compare to Figure 1), showing different points where provenance and integrity of software could be established (PI, PE, PB, PD, PS, PA) and Verified (VI, VE, VB, VD, VS, VA) using software signatures. Abbreviations are defined in Table III.
TABLE III: Different Signing use Points as noted in  Figure 4 and subjects who employ software signing at these points. We mark subjects who consider them optional for reasons of demand by customers, nature of products, etc.
Point Description Subjects
SOURCE CODE
PI Internal code contributors sign their commits S2-S11, S14-S17 (S4, S8, S14 – Optional)
PE External code contributors sign their commits S7, S10 (S11, S14 — Optional)
VI Verification of signatures from Internal contributors S7, S14, S17
VE Verification of signatures from external contributors
PSC Signing after code reviews/audits S2-S11, S14-S17 (S4, S8, S14 — Optional)
BUILD
VSC Verification of source code signatures before build S2, S3, S7, S14
PB Signing of build output S1-S5, S7, S8, S11-S15, S17 (S8 —Optional)
DEPLOYMENT/PACKAGE/ARTIFACT
VB Verification of build signature before Deployment or final artifact S7, S14
PA Signing of Final Software product All S1–S18
THIRD-PARTY DEPENDENCY
PD Signing after verification and certification of dependencies S6, S7, S9, S16, S17, S18
VD Verification of Signatures from external dependencies S9, S11, S17 (S9, S11, S17 — Optional)
VA Verification by final customers Difficult to determine
Using our initial analysis plus the recommendations and philosophy of several frameworks, which recommend a continuous verification of integrity and trust in code alone (not humans) [20, 92], we present a refined form of the software supply chain factory model as shown in Figure 4. Our model comprehensively identifies points in the software delivery process where the security property of provenance is expected to be established or verified. We compare this with subjects’ reported usage of other security methods for these factory model components, and second with framework (and standards) prescriptions.
5.1.1 Source code
First, we ascertained whether our subjects used any form of external code contributors in their software creation process. Out of 10 subjects who reported a form of collaboration with open-source or other external contributors, only 5 subjects used external code contributions in the production code of projects they owned (both commercial and OSS). We then proceeded to determine if any form of signing is used to establish provenance for points PE, PI, VE, VI, and PSC.
External contributors sign commits (PE): Signing is mostly required by teams who use external code contributors. Out of 5 subjects whose teams used external code contributions in their production code, 4 of these subjects require that commits be signed. However, two of those do not require signing for open-source projects or noncritical projects. Only 2 of our subjects unconditionally require signing from their external contributors. As S2 put it, “I got something signed from somebody I don’t know…So we do the code review, and then we have a sign-off by somebody on our team that signs off from that code”.
Internal contributors sign commits (PI): Signing is almost always required by teams for their internal code contributors. 14 of 18 subjects report that their team required some sort of commit signing from internal contributors. Subjects who list this as an optional requirement (3) cite similar reasons as in PE above.
Verification of signatures from contributors (VE&VI): While signing is relatively required, it is not verified as a check for security and provenance of code contributions from contributors. Techniques like code reviews (and security audits), security/vulnerability scans, tests, etc, are the most used security methods for assessing the security of code contributions. S10 opined to this, “I think everybody on the team does sign their commits, but we don’t really verify it…If somebody stopped signing their commits, I don’t think much would really happen”
Code is signed after reviews/audit (PSC): Signing here is common similar to PI. Subjects who report that signing is required in PI, also report that all activities related to code reviews, or audits are (cryptographically) signed off on.
5.1.2 Build
Next, we extract from our subjects’ responses how signing is used in the build process of software products. For most of the subjects, the build and build signing (for subjects who do so) are automated. Our model expects that the source code to be built should be verified (VSC), and then when the build is successful we expect another round of provenance certification (PB).
Verification of source code signatures before build (VB): Signatures from source codes are rarely verified before the commencement of the build process. Two subjects mentioned some sort of rigorous verification at this point. Alternate means of security at this stage are similar to the security techniques reported for the build process.
Signing of build output (PB): 13 subjects reported the use of signing at this stage of the software delivery process. 10 subjects reported signing is used alongside attestations, SBOMS, and other forms of metadata.
5.1.3 Package/Artifact
Next, we analyze the use of signing at the final stage of deploying software products.
Verification of build signature before deployment of final artifact (VB): Similar to the other verification steps, signatures from the build phase are rarely verified before the deployment. Only two subjects mentioned verification at this point. For example, S7 states, “We sign our code and then we also verify that it was built in our specific build system”.
Other subjects mentioned that the security of the entire build process comprises techniques like an automated build pipeline with Workflow definitions, use of build policies, and Reproducible builds. These are complemented by code scanning tools as in the case of the source code stage.
Signing of Final Software product (PA): All 18 subjects said that signing is done at the deployment stage to establish provenance for the organization on each of their products.
5.1.4 Third-Party Dependencies
For third-party dependencies in the Software Supply Chain, we evaluated how software signing is used as a factor influencing dependency selection, and how software signing is used to verify the authenticity of dependencies in practice.
Presence of a Signature as a factor for Dependency Selection: For most of our subjects, signatures on a third-party dependency do not affect their choice in selecting such a dependency. Only 3 of our subjects like to see this, treating signatures as a signal that the owner is security conscious. Other security factors that influence the selection of third-party dependencies as mentioned are: the use of the OpenSSF’s scorecard tool, the dependency’s known vulnerabilities, level of activity, licensing, and age.
Verification of Signatures from external dependencies (VD): Signatures on third-party dependencies are rarely verified in practice. Only three subjects reported some form of verifying signatures from third-party dependencies. As S11 said, “If it is open source, then probably no…But if it is commercial off the shelf, we want to make sure those are signed.” However, most of our subjects reported several methods for assessing the authenticity of a third-party dependency. These methods include code and vulnerability scanning (e.g., Coverity), source rebuilding, internal mirroring of dependencies, and reliance on ecosystem internal checks (e.g., Go package manager’s checks). Other subjects perform internal reviews of dependencies, while some have a centralized dependency selection and management process, where dependencies are hosted and maintained by the organization. A few subjects noted that their organization discourages or forbids third-party dependencies.
Signing after verification and certification of dependencies(PD): A few subjects reported that third-party dependencies are thoroughly reviewed using one of the aforementioned techniques before being signed off for use by a team member. Six of our subjects indicated that this process is followed in their software production workflow. For example, S17 described their team’s process like this: “We have a signed report that’s signed by us, and we trust us…that says, ‘Hey, I scanned this third-party dependency, did all these things. I don’t have a signature for that, but I ran all these extra checks and then signed those checks. So now we have a record”’. In contrast to that attestation approach, S18’s team ingests third-party dependencies and signs the internal versions directly.
5.2 RQ2: Challenges affecting the implementation of software signing in practice
Finding 2. Subjects identified several common themes as challenges faced in practice as obstacles to effectively implementing software signing. We categorize these broadly; 1. Technical Challenges: These include – key Management and Infrastructure issues, compatibility issues with existing systems and different architectures, and lack of signature verification 2. Organizational and human Challenges: These include – expertise in setting up Signing implementation, operationalization of the signing procedure, developer attitude, and Resources to set up the signing process
Our next research objective was to understand the various challenges practitioners face while implementing software signing in practice. We grouped subjects’ responses under three main themes for challenges faced by practitioners in practice; Technical, Organizational, and Human factors. Table IV gives a summary. Due to space limitations, we focus on only a subset of the individual challenges.
TABLE IV: Challenges to software signing Implementation in practice. We have categorized related challenges highlighted by subjects. Two uncommon Technical challenges are described in Appendix E.
Observed Challenges #Subjects(#Orgs) Proposed Solutions
TECHNICAL
Key management 10 (9) Use of Keyless Signing (e.g., Sigstore)
Compatibility issues 6 (6)
Lack of Verification of Signatures 6 (5) Signed Metadata, Component Data Management
Ease of use/Usability 4 (4) Usable signing tools (e.g., Sigstore), Documentation
No unifying Standard 2 (2)
Lack of Verification of Signatures 6 (5) Signed Metadata, Component Data Management
ORGANIZATIONAL
Operationalization of the Signing process 4 (4) Automating Signing
Resource to Set up Signing 3 (3)
Creating Effective Signing Policy 2 (2) Regular Process Feedback mechanisms
No Management incentive to Sign 2 (2)
Bureaucracy 1 (1)
HUMAN
Expertise in setting setup and use of signing 5 (4)
Developer Attitude to Signing 3 (3) Automating Signing
Lack of Demand from Customers 1 (1)
5.2.1 Technical Challenges
Key management: Key management Issues were the most reported issues. Issues under this category vary from key management, key distribution, key generation, and public key infrastructure (PKI) issues.
Key management issues are the highest categories recorded. S13 articulates this, especially concerning key distribution: “a lot of the ways in which I think previous teams’ software signing hasn’t been useful is…key management…in terms of how the private key is managed but also in how the public key is made available.”
Some subjects discussed recent signing technologies that implement keyless signing (e.g., Sigstore) as a possible solution. S18 said: “Sigstore comes in…keyless signing…you don’t have to worry about long-lasting static SSH keys, or keys being compromised, or rotating keys…you don’t have these big servers to store all these signatures”. S17 also noted, “[Sigstore’s] other strength…I can associate my identity with an OIDC identity as opposed to necessarily needing to generate a key and keep track of that key and yada, yada. So that’s super useful because I could say, ‘Oh, this is signed with [a] GitHub identity.’ So unless my GitHub identity has been compromised, that’s much better.”.
Compatibility issues: Compatibility issues arise during the integration of software signing systems and tools with other tools within the software engineering process, including other security tools and the artifacts to be signed.
S2, S7, S17, and S18 highlight issues surrounding the compatibility of newer signing systems with existing software systems in the software engineering process. As S18 put it, “There are teams who’ve been doing this for quite a while…Selling that is a big challenge, right? Because we need to prove that it is much more secure than what has been there in place.”
S8, whose organization’s software product is built for different platforms and hardware architectures, noted compatibility issues across various architectures: “Every platform evolved separately…macOS…Windows…The requirements are changing and evolving independently”.
Compatibility with other security tools can pose challenges too, as noted by S5: “The vulnerability scan report [is for]…code…CVE published…[but] signatures, I don’t think it will pick those things up.”
Ease of use/Usability: Usability of the signing tool or implementation was a common challenge hampering effective use in practice. This was often expressed as developer experience, e.g., S17 said: “For a developer, make sure that they can sign the commits easily”. This point was also raised at a team level by S13: “I think that awareness of the value of signing and sophistication around how signing is performed has increased significantly, especially in recent years. But it has also resulted in a lot of smaller teams deciding that this is not a manageable thing.”
With respect to documentation, S11 said: “I think there are multiple challenges there. When you build some signing infrastructure, you want to make it a self-serve so the teams can onboard themselves, how well you have your self-serve playbooks written down, documentation on using those. I think documentation can be improved.”
Lack of Verification of Signatures: Lack of verification of signatures is also quoted by several subjects as a main challenge. This arises from the factors like signing serves as a “checkmark” security tool (S4: “most people who are strictly signing all commits don’t have a meaningful process for verifying those signatures or for enforcing any policy around them. They just do it because”), and lack of customer interest (S14: “We have tried to … get potential customers and customers interested…but as far as I know, none of our customers…verify our signatures on images”).
Verification Of Artifact before Signing: Verification generally was highlighted as a problem, especially verifying inputs to the build process of the software delivery process. The problem noted here is that the content of software artifacts is not readily verified before signing: as S16 put it, “if you start with the component, the component gets put into a subassembly. The subassembly gets put into a package. The package can get put into a broader package or whatever. I think the people are signing. But do you know that the sum of all the pieces are accurate? … what are you really signing, and at what level?”. S11 further noted that different teams could sign arbitrary artifacts due to shared signing infrastructure.
5.2.2 Organizational Challenges
Operationalization of the Signing process: The implementations of software signing are typically neither uniform nor centralized across different organizations. Various teams are largely responsible for crafting their own solutions. As S6 said, “I can’t be more specific, but today, I’ve talked to three different teams with three different processes”. S15 suggested that automation could improve uniformity: “We try to break down the barriers with really good examples, by having common configuration of CI through similar GitHub workflows, and stuff like that. So it’s easier for people to use.”
Resource to Set up Signing: The issue of organizational resources to set up efficient signing tooling across an organization was also another cause for concern for some of our subjects. Subjects discussed this in terms of finances, time, and engineering effort. S14 said, “We need to do a lot of signing…hundreds of container images [daily]…It was a pretty significant engineering effort to build a pipeline that can handle all those signatures.”
Creating Effective Signing Policy: Policies surrounding the implementation of software signing by teams are generally not considered highly effective. These policies include elements such as role-based access, key usage, and key issuance. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S3: “… the other thing is policies, creating a policy that is, for one, it’s hard to do. And then the other part is creating an effective policy and knowing that it’s effective”
A suggested solution to address this issue is the use of regular feedback mechanisms to assess the security policies and processes of different teams within the organization. As S11 said, “We want to be more proactive…internal audits, and an assessment methodology like the NIST, I think CISF.”
No Management incentive to Sign: A lack of management incentives was also a challenge. S11 said, “The investment, the time that the engineer is putting into…signing…before a release, does it get the right attention from the senior management? Do they view it as a time well spent for the product?”
5.2.3 Human Challenges
Expertise: Some subjects noted that setting up signing tools depends on the engineer’s expertise. Unfortunately, subjects felt that most engineers lack the necessary experience and skills for this task. As S9 said, “there’s a lift there as far as…expertise”.
Developer Attitude to Signing. Software developers’ attitudes toward software signing pose challenges to the signing process. Disinterest and resistance to adopting new tools are the primary attitude-related obstacles cited. S10 observed: “Most developers don’t really understand this stuff and don’t really want to think about it…[they are] happy if they…see a big green tick[mark]”. S6 added: “They refuse to get off their old tools”
Lack of Demand from Customers: Most customer-driven organizations report that software signing is challenging to implement because customers rarely request this feature in their requirements. S14 put it like this: “Sign[ing] our images was there from day…we haven’t found many customers that are too excited about the signing aspect”.
5.3 RQ3: Perceived importance of Software Signing in mitigating risks
Finding 3. Subjects disagreed on this point. Some view signing as crucial, while others treat it as a secondary consideration in their overall strategy.
Here, we aimed to understand how important practitioners perceived software signing to be in their software delivery security. Opinions vary on the significance of signing in the software delivery process.
5.3.1 As Crucial to Provenance and Integrity
Many of our subjects considered provenance an integral part of security, and felt that software signing is (or will be) a guarantee of that. S4 said, “I think signing is massively important. Signing for open source is probably the single most important thing for software supply chain security”.
5.3.2 As a Secondary Security technique
Some subjects believe that software signing is a supplementary method used to ensure the integrity of primary security techniques. They often deemed it important only in conjunction with methods like Attestations and other Manifest forms such as the SBOM. As S2 said, “I think the most important thing is accurate metadata collection and centralization. And even if that data’s not signed, we can still get a lot of usefulness out of it. Signatures, in my opinion, they offer an additional layer of security on top of what we should be doing, which is metadata collection”. S17 similarly views signing as a final check on top of other security practices: “Signing was essentially the way of saying, ‘I have checked, we did all the [security practices]’.”.
5.4 RQ4: Impact of Supply Chain Security Awareness on Software Signing Practices
Finding 4. 1. Software supply chain failures usually require immediate fixes to the malfunctioning or compromised components. Large-scale failures typically trigger a comprehensive reassessment, leading to changes in the security strategy of a team or organization. 2. Security regulations, frameworks, and standards drive the adoption of signing and other security techniques, but some subjects criticized these obligations.
In this section, we explore how experience with software supply chain security influences a company’s decision to adopt and implement software signing. We asked subjects about their personal and organizational experiences with software supply chain failures (attacks, vulnerabilities, and other incidents) and how these experiences affected their team’s or organization’s software supply chain security decisions, particularly regarding software signing. Additionally, we queried subjects on the impact of software supply chain security frameworks, standards, and regulations on their use of software signing and other security techniques.
5.4.1 How experiences with SSC attacks influence the use of software signing
TABLE V: Experienced software supply chain failures and type of Security change triggered.
Type # Subjects (# Orgs) # Direct fix # $\Delta$Signing # $\Delta$General Security process
Attack 4 (4) 4 2
Vulnerability 9 (6) 9 1 2
Other Incidents 5 (5) 5 2
Software supply chain security failures typically have direct fixes, except in cases of severe attacks. While only 4 of our subjects indicated experiencing software supply chain attacks, only one of these attacks (where the subject was hired to help fix the aftermath) was highlighted as major, resulting in collateral damage and leading to a complete overhaul of the software and security processes. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S4: “I would say that at this point it’s probably difficult to find any process at that organization that wasn’t inspected and potentially changed as a result of that. Every aspect of how everything was built there was changed … There was no choice in that situation.”
Some subjects report that software supply chain security failures led to a broader change in their team’s or organization’s security posture. For example, S18 describes their team’s response to the Log4j vulnerability, which was present in their system. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S18: “… it was a reality check to know that even companies like ours … can be attacked through open source libraries, … Asset inventory is the biggest thing that is lacking where you don’t know what is currently running in your environment. And if it is running, you don’t know how old that is, or what are the dependencies which is currently running in production. So that pushed us towards making sure that we have a continuous inventory of container images, which are currently running, for example, for Kubernetes, and then there are initiatives like GUAC.”
Generally, most software supply chain reported by our subjects were not severe, and as such most of them received direct fixes. This could be swapping out a defaulting tool or dependency, re-releasing an affected release, publishing security advisories, etc. We highlight some of those responses. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S5: “So what happens is that, if the current version where that security risk is present, if that is also publicly available, we have to create a security advisory to let everyone know that if you are working with this version of our software, you should upgrade to the next version which contains the fix. And we let them know what the fix is and if they are affected by the fix.” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S1: “to my knowledge it did not cause a major restructuring or change that is … Yeah, I would say that they probably patched any of the security issues that they had.”
The only instance where a signing-related fix directly resulted from a software supply chain security failure was when the subject team’s product was a safety-critical application. This team added intoto’s signed attestation capability. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S2: “we started reaching for the In-toto project, and we saw that it allowed you to decouple the security metadata collection and validation. You can abstract that away from the actual build process itself. So that decoupling allowed us to solve this problem of how do we know that this opaque process for the software, that opaque certification, how do we know that was followed without actually knowing the details of what that process is?”
5.4.2 Influence of security regulations, frameworks, and standards
TABLE VI: Regulatory and Standard Influences on Security Technique Adoption: Analysis of Subjects’ Responses. 4 Subjects say they are not influenced by either.
ANY INFLUENCE?
#Subjects (Yes) #Subjects (None)
Regulation(EO-14028) 11 7
Other Standards 8 10
The adoption of software signing and other security techniques is heavily influenced by regulations and other standards. Specifically, we asked our subjects, “Are these strategies (software signing implementation & other complementary security methods(earlier listed by subjects)) influenced by regulations and standards?”. While answering generally, most of our subjects agreed that regulations, particularly the E0-14028, played a key role in the way their team and organization have implemented software supply chain security. Although software signing is not explicitly mentioned, Some subjects did note that their use of SBOMs was directly a result of regulations or standards. We only had 1 subject who noted that the standards influenced their use of signing in certifying attestations(SBOMs).
S1 highlights this influence of the executive order: “Yeah, I would say certainly the collection of SBOMs and doing that more broadly has been a response to things like the executive order, at least for our team specifically. I would say, yeah, the executive order played just a strong role in what we are starting to do now. Whether or not some teams are now trying to be preventative, that’s a good question.” S7 also adds to S1’s point: “we keep the executive orders in mind when we build our product…”
Some of our subjects (3) only relied on standards (not on regulations) for security recommendations. For example: {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S18: “… as an organization, currently we don’t have any government contracts. And I think the executive order is mostly targeted towards companies, organizations who are having contracts with the government and they need to produce certain SBOMs and things like those. So our push towards supply chain or better security was not influenced by the executive order.”
S18 also adds: “But I think in terms of where we are at, I think in terms of frameworks, as you mentioned, we do look at some of these frameworks like SLSA,… There are quite a few …, and each of them target different parts of the supply chain. But as an organization, it’s really difficult to keep track and follow each of these standards of frameworks, to be honest, because we have limited resources, we can’t just keep on handling those. So we are looking at SLSA, or at least trying to get closer to where we should be for the SLSA level. Again, it’s not written in stone that we need to have level one, level two, level three, things like that, but that’s something which we strive to achieve sooner.”
S14 adds to this line of thought: “The recent interests from the US federal government and software supply chain security, SBOMs, and the executive order from 2021, it’s hard to detect very much influence on our security because of those … I think when observers of software security read that executive order and then commentary in the aftermath, they often think that SBOMs have become mandatory…”
5.4.3 Criticism of Regulations and Standards
Also, our subjects identified several issues with the current set of software supply chain security standards and regulations that hinder their influence on the adoption of software signing and other security tools. We highlight four kinds of issues.
(a) Regulations don’t specify standard data exchange format for security tools: Criticism of standards and regulations here lies in their inadequacies to specify or promote standard data interchange formats. This is highlighted by the lack of data formats for certain current security techniques such as Software Bill of Materials (SBOMs), other manifest-type techniques, signing tools, etc. The absence of standards for these techniques leads to numerous undesired practices, such as the use of vulnerability scan reports as an input for SBOM data. This ultimately leads to compatibility problems among different security techniques.
{quoting}
[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S16: “…one of the gaps with the executive order is, yeah, you give an SBOM now. But what are they going to do with it? Nobody knows yet. And how do you compare it? Because even right now, going back to the tools, there are different versions of SPDX. There are different versions of Cyclone DX. They’re the two big ones. This goes back to tools. There are a bunch of tools out there that will convert say 2.0 back to three and you come up with different answers using different tools.” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S3: “SBOMs are a good example of something that has a standard, but all of the implementations end up being different enough that they’re very difficult to translate from one format to the other or even get consistency within the same format type.”
(b) Regulation and standards are just checkbox rules:
Here, subjects argue that standards and regulations function more as compliance checks rather than as security enhancers. They point out that these standards and regulations lack security enhancement features, such as implementation models. Instead, they primarily promote compliance-type features, such as guidelines and best practices lists. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S17: “So there is currently a disconnect between regulations and security. So compliance and regulations keep people out of jail, keep people from getting fined, [while] security stops you from getting hacked. And the problem there is sometimes those two things can conflict. And compliance is there to make sure that everybody is following reasonable sets of rules and security is there to actually implement that and create the software and the processes that hit that … Regulations help provide the requirement. They don’t necessarily make security better on their own. And there’s always going to be a balance because what compliance tells you is compliance helps you with making sure that everybody’s following the standard … the only way for me to be compliant to this is to ignore the actual problem and just check the box.”
(c) Difficult to implement in the SDLC: A group of participants highlighted the challenges encountered when attempting to integrate the recommendations of security standards into their software engineering processes. These challenges primarily arise from the difficulties in training engineers on the application of these standards, and the deficiencies in the technical implementation of the standards themselves. S12 makes this point clear in their response; {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S12: “… just having the executive order was just another way to explain it to the higher-ups… I don’t think that the executive order in particular affected the business … There’s kind of milestones. You don’t need to know exactly what the ins and outs are technically of achieving them … But it’s a lot of work getting that in place and upskilling all of your existing developers, et cetera. So yeah, anything that requires going to all 4,000, 10,000 of your developers in some of these massive organizations and making sure every single one of them does something different, it is a lot of work.”
(d) Standards are not formalized: Criticism regarding the formalization of standards, as expressed by some participants, reflects the lack of consensus surrounding these standards. Participants argue that since these standards are not necessarily created by a majority of industry consortia, they should not be referred to as ‘standards’. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S13: “the term standards is imprecise. So some people think of standards as being very formalized but some people would say any documented practice is a standard. So there’s the SLSA project in it, and one might argue that that is a standard or one might argue that it’s not a standard, but it is a documented set of recommendations. And if people are implementing those, you could argue that they’re implementing a standard.”
6 Discussion
6.1 Recommendations from this study
Improving Signature & Artifacts Verification in Signing workflows.   Our refined software supply chain factory model analysis has highlighted a significant gap in the current practice of signing - signature verification. This issue was underscored by several subjects in our study.
In certain automated workflows, the act of signing is often treated as a mere formality, serving as a “checkbox” that, once done, allows engineers to carry out their usual tasks (e.g., creating pull requests or merging code into the main branch). Thus, this checkbox provides very little security properties, and requires effort to provide meaningful security gains.
Further, while artifact verification, in this context, refers to conducting a thorough check of the integrity of the software artifact by means of identifying the signer. However, the use of signed attestations (i.e., are statements about a particular action or property in the supply chain) can further verify the authenticity and integrity of the process and will become crucial to further improve the security posture of organizations.
Future works in this area should focus on proposing improved tooling around existing or new software signing tools to include these capabilities in their design. This enhancement would not only improve transparency but also reduce the costs and expertise requirements for setting up such infrastructure.
By addressing these issues, we can ensure that signing is not just a procedural element, and instead plays a crucial role in maintaining the integrity and authenticity of the software development process.
Addressing Incentive Challenges in Software Signing.  One of the most frequently reported challenges in the realm of software signing is the lack of proper incentives for its adoption. These incentives can range from support from management, the ease of using signing tools, as well as comprehensive education on the importance and application of signing.
While some of these issues, such as the need for educating practitioners on the significance and application of signing, can be addressed with relative ease, others necessitate more elaborate planning. For instance, fostering management support might require demonstrating the tangible benefits of software signing in terms of enhanced security and reduced risk of software tampering.
Moreover, though the user-friendlness of signing software has been evidenced by various “why Johnny can’t” papers, more research is needed in lowering the adoption bar for software signing systems.
Regardless of the complexity of these challenges, investing in these areas is paramount for enhancing an organization’s security posture. By fostering a culture that values and understands the importance of software signing, organizations can ensure the integrity of their software and protect against potential security threats.
Future research could focus on developing strategies to address these incentive-related challenges and exploring their effectiveness in different organizational contexts. This could provide valuable insights for organizations striving to improve their software signing practices.
Criticism of Current Standards and Regulations.  A significant finding of our study is the criticism directed towards current standards and regulations, particularly concerning their creation process and the difficulty of their implementation. These issues could be mitigated by ensuring that these standards and regulations are informed by a broader section of the industry consortium. This would involve gathering input from a diverse range of practitioners, including those from different sectors, roles, and levels of expertise.
While frameworks like SLSA already involve collaboration among multiple organizations, we advocate for a more inclusive effort. This would target a larger group of practitioners, ensuring that the standards and regulations are not only comprehensive but also practical and implementable across various contexts.
In the future, research could focus on exploring effective strategies for fostering such inclusive collaborations and assessing their impact on the quality and applicability of the resulting standards and regulations. By addressing these issues, we can work towards standards and regulations that are not only robust and comprehensive but also widely accepted and easily implementable. This would significantly enhance the security and integrity of software signing practices across the industry.
6.2 Contrast to prior work
Comparisons with Previous Quantitative Studies on Software Signing.  We compare our findings with Schorlemmer et al.’s [5] quantitative study on software signing in open-source ecosystems. Their results indicate that major software supply chain failures (e.g., SolarWinds, Codecov attacks) and dedicated tooling had no significant impact on the quantity of signatures. In contrast, our findings for the non-open source industry show the opposite. This difference is expected due to the substantial incentives, such as commercial and reputation losses, that drive greater adoption in the non-open source industry. Additionally, our study supports this finding, as some subjects noted that contributions to the open-source ecosystem often involve relaxed security requirements.
Distinctions of Refined Model Compared to Existing Frameworks.  While all software supply chain security frameworks agree on the defined threat model, our refinement of this model differs from existing frameworks in several ways.
  1. 1.
    SLSA: The SLSA framework primarily focuses on securing the build process within the software engineering lifecycle, offering recommendations as a series of incremental security hardening levels. In contrast, our refinement addresses the entire software supply chain, not just the build step.
  2. 2.
    SCIM and S2C2F: This framework provides threat-reduction recommendations without specific implementation details. Our approach, however, emphasizes the continuous establishment and verification of provenance across the software supply chain.
  3. 3.
    Hammi et al. [93]: While their work proposes a blockchain-based cryptographic tool, we do not advocate for a specific tooling solution. Instead, our model is designed for industry and software teams to create an implementation workflow that is independent of specific tools. Additionally, our findings are supported by insights from industry practitioners, enhancing the practical applicability of our model.
7 Conclusion
Signature-based supply chain provenance is a topic of increasing importance, driven by cyberattacks, executive orders, and new industry standards. Although the literature contains many studies of signing as a technology, there have been few studies of the use of signing in engineering contexts. Those that exist have focused on open-source, or on the mechanics of signing. In this paper, we present the first qualitative examination of signing in engineering organizations’ security strategies. We interviewed 18 high-ranking practitioners across 13 organizations, including C-suite staff, principal and senior engineers, managers, and consultants. Our four main findings are that (1) Software products are being signed, but the signatures are not yet being validated; (2) The factors that hamper signing include infrastructure and compatibility issues, lack of expertise, developer and organizational attitudes, and lack of resources; (3) Signing is considered an important guarantee of provenance, but it is not the main technique these organizations consider in their cybersecurity postures; and (4) Organizational behaviors are being influenced by major failures (e.g., SolarWinds) and by regulations and industry standards, but some of our subjects expressed concern about the development process and quality of those recommendations. Our results indicate many opportunities for technical and operational advances in the use of software signing to improve software supply chain provenance.
8 Replicability
For anonymity reasons, we do not plan to share the raw transcripts of our interviews. However, our interview protocol, codebook, etc. are available in the Appendices.
9 Acknowledgments
We thank the study participants for contributing their time to this work. Taylor Schorlemmer and Wenxin Jiang assisted in typesetting. We acknowledge support from Google, Cisco, and NSF #2229703.
Appendix A Outline of Appendices
The appendix contains the following material:
  • Appendix B: The interview protocol.
  • Appendix C: The codebooks used in our analysis, with illustrative quotes mapped to each code.
  • Appendix D: Analysis of saturation in our data.
  • Appendix E: Uncommon challenges, omitted from the main paper due to space limitations.
Appendix B Interview Protocol
Table I gave a summary of the interview protocol. Here we describe the full protocol. We indicate the structured questions (asked of all users) and examples of follow-up questions posed in the semi-structured style. Given the nature of a semi-structured interview, the questions may not be asked exactly as written or in the same sequence but may be adjusted depending on the flow of the conversation.
We include Section D of the protocol for completeness, although it was not analyzed in this study.
TABLE VII: Our full interview protocol. The table also indicates the history of the protocol development: removed (strikethrough), changed ($\star$) and re-arranged ($\dagger$) questions.
Themes Sub-themes Questions
A: Demographic A-1: What best describes your role in your team? (Security engineer, Infrastructure, software engineer, etc)
A-2: What is your seniority level? How many years of experience?
A-3: What is the team size?
A-4: $\dagger$ What are the team’s major software products/artifacts?
A-5: What type of Organization/company?
B: Software Supply Chain
Failure Experienced
B-1: Describe briefly the team’s process from project conception to product release and maintenance – This is to understand the unique context of each practitioner’s software production case.
B-2: $\star$ What do you consider a Software Supply chain attack/incident to be?
B-3: Can you describe any specific software supply chain risks (or incidences with third-party dependencies, code contributors, open source, etc.) that your team has encountered during your software development process? How were these addressed? (Alternatively – How did this affect the decision to implement software signing?
[…]
B-5: What is your team’s greatest source of software supply chain security risk? Project components(Third-party dependencies, build process, code contributors, etc) $\dagger$ Between the Software Engineering Process vs project components which constitute a greater source of risks?
C: Software Signing in Mitigating Software Supply Chain Risks C-1: If no incident has been recorded (depending on the answer from  LABEL:sec:_B-sec-experienced): why did the team choose to implement software signing? Is software signing the team’s major strategy to secure its supply chain? Any other complimentary security efforts and methods? Are these strategies (software signing and complementary security techniques) influenced by regulations and standards? What challenges or obstacles did your team face while integrating Software Signing into your supply chain security practices? $\star$ How do different team members contribute to the implementation of Software Signing? What roles and responsibilities are involved? (After Practise). Possible Follow up – Do you think Software Signing on its merit (if implemented right) is good enough to completely secure a supply chain? (Added after Pilot).
Third-party dependencies C-2: What are the team’s peculiar selection strategies for Third-party dependencies? How does the presence of a signature influence this decision? How is the authenticity of this signature verified? Any other methods/practices to ensure the trustworthiness of the third-party dependencies before integrating them into your projects? Possible Follow-up – Can you describe these methods/practices?
C-3: What influences the team to discontinue the use of a package? How does the Signature on the dependency influence this decision?
C-4: How does the team manage its third-party dependencies’ security? How do you maintain awareness of potential vulnerabilities or threats related to third-party components?
First-party source code C-5: Is software signing a requirement for team developers (insider threats) and open-source contributors (if any)? how does the team use software signing to protect its source code (what parts of the process is signing required e.g. Commit signing)?
Build Process C-6: (How) Does the team utilize signing in its build process?
Package Artifact/Deployment C-7: How is Signing used to protect the following from compromise? how Artifact’s build binaries/deployment pipeline Artifact’s repository
C-8: $\star$ How does the team evaluate the effectiveness of their Security processes/Signing Implementations? Any Feedback mechanism?
D: Signing Tool Selection D-1: What software signing tooling does the team use (If not mentioned yet)
D-2: What factors did the team consider before adopting this tooling over others?
D-3: What was the team’s previous signing practice before the introduction of current tooling?
D-4: How does your team implement this tooling (which components of tool does the team use)?
D-5: Have you encountered any challenge(s) using this tool of choice? How have you coped with this limitation(s)? Are there any areas where you believe further enhancements could be made in your current implementation of these methods?
D-6: Have you/your team considered switching this tool for another? AWhat other tools have been considered or are currently being considered?
Appendix C Codebooks
Table VIII describes our codebook, with code, definition, and example quote tagged with that code.
TABLE VIII: Excerpts from the final codebook we used for our analysis. Sample codes under each theme give examples. Note that the theme, “Others and Specific tooling” is not discussed in the paper.
Code Definition Sample Quote
BUILD
Build Infrastructure of team Nature of Subject’s build process or build Infrastructure. Example- when subject mentions that their build or CICD is automated S7: "Our build system is typically GitHub actions."
SOURCE CODE
Signing Implementation in Source Code (External Contributors) Subject’s manner of implementing Software signing in first-party source code produced by external contributors like 3rd party vendors, example, commits by contributors are signed S2: "I don’t know who these people are … I got something signed from somebody I don’t know … So what we do though is that we do the code review, and then we have a sign-off by somebody on our team that signs off from that code"
DEPLOYMENT/PACKAGE/ARTIFACT
Signing Implementation in Package/Artifacts Subject’s manner of implementing Software signing in Deployment of Product or packaging of created artifact (Release) S10: "git commit signing on the code and then signatures on the final artifact. "
THIRD-PARTY DEPENDENCY
Signing Implementation in Third Party Dependencies and Tools Contains code category related to 1. Influence of Signing in Selection of third party dependencies and 2. Influence of Signing in discontinuation of dependency use S13: "So most software package ecosystems that I consume packages from do not support signing consistently enough to put much stock in that. "
EFFECT OF SSC FAILURE ON ADOPTION
Experienced Software Supply Chain Attack Subject fell victim to a malicious orchestrated software supply chain attack. For example, an organization falling for phishing S4: "The build system itself was compromised by a state actor…somebody compromised the system and lurked for a long time and built very bespoke malware, specifically tailored to Orion, to mimic Orion’s network traffic, all of that other kind of stuff…NSA and GHGQ both said that this was Russian foreign intelligence that did this to SolarWinds."
SPECIFIC TOOLING
Sigstore Criticisms Criticism and weakness associated with implementing SigStore for signing S3: "we do have some concerns about the way Fulcio operates as its own certificate authority. So we’ve been looking at things like OpenPubkey…removing that intermediary…to do identity-based signing directly against the OIDC provider."
PERCEIVED IMPORTANCE OF SIGNING
Provenance and Integrity Subjects mention that their current or previous teams use software signing in a major role for establishing provenance and integrity or other similar roles. S4: "A few years from now, everybody expects provenance…they don’t install anything if they can’t prove where it came from…I think signing is massively important. Signing for open source is probably the single most important thing for software supply chain security"
SSC STANDARDS AND REGULATION INFLUENCE
Influence of Regulations and Standards on SSC Security Strategies How regulations and standards influence their SSC security strategies on signing, or use of other security methods S8: "My external looking in perspective of things, is that there’s a mix of anticipation of regulation, as well as reaction to regulation. For the most part, there’s a lot of trying to predict what is going to come, what the government is seeking for as far as tech regulation, and trying to get ahead of that curve."
CHALLENGES
Challenges experienced in the use of Software Signing These are responses from subjects when asked about challenges to them implementing Software Signing generally and not just criticism of a signing implementation like Sigstore/GPG/Notary etc. For example when the subject talks about Key management, Bureaucracy, motivation S7: "I’m not confident that those customers actually ever verify those signatures."
OTHERS
Choice of tool - General Reasons What influences choice of software supply chain security tools except signing tools S17: "A lot of the tools are not particularly easy to use "
Appendix D Saturation Analysis
Figure 5 shows the saturation curve associated with our interviews.
Figure 5: Saturation curve. Each interview covered many topics in detail (orange line, blue line). However, as the green line shows, our results saturated (i.e., stopped observing new perspectives) around interview 11.
Appendix E Uncommon Challenges
E.0.1 Managing Signed Artifacts
Managing signed software artifacts post-signing should follow good practices, but this is currently a challenge. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S9: “Also managing all the artifacts that are signed … It’s great that you sign things, but what do you do with it afterwards? I think that’s kind of something that we’re facing right now, especially when it comes to container management. Great, we signed containers, but now what? … Okay, great, now what? All those downstream things that we’re trying to figure out.”
E.0.2 Identity Management/Authentication
Managing and authenticating the identities of team members authorized to sign at various stages of the software delivery process poses challenges in some instances.
{quoting}
[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S1: “in general one of the challenges that I hear from other security architects and security leads, and program managers, and people at that level, is key management and identity management as being a barrier for signing and security and authentication in general.”
Possible solutions suggested and used by some participants include the use of short-lived or constant key rotation. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S7: “We really rely on short-lived tokens for authentication. I think that is one way that we really try to restrict access to our system and make sure that there’s no long-lived secrets, there’s no long-lived access tokens anywhere in our system. For any signing that we do, I think those keys are rotated quite frequently. So, that’s another thing we do to prevent any attack should key information be leaked. I’m trying to think what else.”